OpenStudy (anonymous):

Standard form of a quadratic function. y=a(x-h) square +k 1. y=1/2(x+3) square +2 how would you graph this ?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Start with the graph of y = x^2 Compress it vertically by a factor of 2 to get y = 1/2x^2 Then shift it over 3 units to the left to get y = 1/2(x+3)^2 Finally, shift the entire graph up 2 units to get y = 1/2(x+3)^2 + 2

jimthompson5910 (jim_thompson5910):

Alternatively, you can plug in values of x to get corresponding values of y. This will give you a series of points which you can then plot and draw a curve through to graph y = 1/2(x+3)^2 + 2

OpenStudy (anonymous):

wait i dont get it really can you show by doing it like f(-2)=1/2(-2+3) square+2 and then try to work it out and i choose -2 because after i plotted -3 i looked and i saw i could use -2 or -4 but i dont know how to work it out

jimthompson5910 (jim_thompson5910):

y = 1/2(x+3)^2 + 2 y = 1/2(-1+3)^2 + 2 .. Plug in x = -1 y = 4 So when x = -1, y is y = 4 So one point is (-1, 4) ----------------------------------- y = 1/2(x+3)^2 + 2 y = 1/2(0+3)^2 + 2 .. Plug in x = 0 y = 6.5 So when x = 0, y is y = 6.5 And another one point is (0, 6.5) ----------------------------------- y = 1/2(x+3)^2 + 2 y = 1/2(1+3)^2 + 2 .. Plug in x = 1 y = 10 So when x = 1, y is y = 10 And another one point is (1, 10) ----------------------------------- Keep going to generate more points

OpenStudy (anonymous):

i still dont get it but can you help me with this one as well, y=2x square

jimthompson5910 (jim_thompson5910):

Do the same: plug in values of x to get corresponding values of y to get a series of points.

jimthompson5910 (jim_thompson5910):

Plot those points and draw a curve through them.

OpenStudy (anonymous):

ok so what would be the vertex and the axis of symmetry for it

jimthompson5910 (jim_thompson5910):

y = 2x^2 is the same as y = 2x^2 + 0x + 0 compare this to y = ax^2 + bx + c What are 'a', b and c?

OpenStudy (anonymous):

a= 2 b=0 c=0

jimthompson5910 (jim_thompson5910):

now plug them into x = -b/(2a)

jimthompson5910 (jim_thompson5910):

to find the x coordinate of the vertex and the axis of symmetry

OpenStudy (anonymous):

oh no see thats a different one from what am doing the formula for this one is y=a(x-h) square +k

OpenStudy (anonymous):

and they just gave me y=2x square which looks confusing

jimthompson5910 (jim_thompson5910):

yeah the different forms can be confusing at first, but you'll get the hang of them

OpenStudy (anonymous):

yea but how would i do that though?

jimthompson5910 (jim_thompson5910):

do what, find the vertex?

OpenStudy (anonymous):

yea and the axis of symmetry

jimthompson5910 (jim_thompson5910):

a = 2, b = 0 plug them into x = -b/(2a)

OpenStudy (anonymous):

would that be zero

jimthompson5910 (jim_thompson5910):

yes, so the axis of symmetry is x = 0

jimthompson5910 (jim_thompson5910):

the x coordinate of the vertex is x = 0

OpenStudy (anonymous):

oh ok thank you, your so my hero :D

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

keep in mind you still need the y coordinate of the vertex to get an ordered pair for the vertex